OpenStudy (anonymous):

find the zeros of the function: 8x^3 +40x^2 +48x

Directrix (directrix):

8x^3 +40x^2 +48x = 0 Factor out the 8x which is a factor of each term. 8x( x^2 + 5x + 6) = 0

Directrix (directrix):

Now, factor ( x^2 + 5x + 6)

jhonyy9 (jhonyy9):

(x*x+2x+3x+6)=0 x(x+2)+3(x+2)=0 (x+2)(x+3)=0 do you can continue it now ?